# LatticeVision: Image to Image Networks for Modeling Non-Stationary Spatial Data

### v1

@@ Abstract
In many scientific and industrial applications, we are given a handful of instances (a ‘small ensemble’) of a spatially distributed quantity (a ‘field’) but would like to acquire many more. For example, a large ensemble of global temperature sensitivity fields from a climate model can help farmers, insurers, and governments plan appropriately. When acquiring more data is prohibitively expensive—as is the case with climate models—statistical emulation offers an efficient alternative for simulating synthetic yet realistic fields. However, parameter inference using maximum likelihood estimation (MLE) is computationally prohibitive, especially for large, non-stationary fields. Thus, many recent works train neural networks to estimate parameters given spatial fields as input, sidestepping MLE completely. In this work we focus on a popular class of parametric, spatially autoregressive (SAR) models. We make a simple yet impactful observation; because the SAR parameters can be arranged on a regular grid, both inputs (spatial fields) and outputs (model parameters) can be viewed as images. Using this insight, we demonstrate that image-to-image (I2I) networks enable faster and more accurate parameter estimation for a class of non-stationary SAR models with unprecedented complexity.
@@ 1 Introduction
Modeling large, gridded spatial data has become a central challenge in many scientific and industrial applications. In situations where data is limited, parametric spatial processes can be fit to the existing data, enabling rapid simulation of additional fields. The bottleneck in this framework is inferring the statistical model parameters using maximum likelihood estimation (MLE), which becomes computationally intractable as dataset size increases [stein2008modeling, sungeostat]. Moreover, spatial data over large domains frequently exhibit non-stationarity, meaning key parameters vary over space, further complicating parameter estimation. Many recent works have replaced MLE with neural networks, mapping spatial fields directly to local parameter estimates [liu2020task, banesh2021fast, zammit2024neural, lenzi2023neural]. Like MLE, these methods are limited to dividing large fields into smaller sections, each assumed to be stationary, and estimating parameters independently for each section. Although significantly faster than MLE, these local, neural estimation methods still scale linearly with the number of sections, and inherently struggle to capture long-range, global context.
@@
Figure 1: An illustration of the main workflow of LatticeVision. Spatial fields are fed into an I2I network, which in turn produces estimates of the non-stationary parameter fields. These are encoded into a SAR model from which synthetic replicates are efficiently simulated.
@@
In this work we introduce LatticeVision, a global estimation and emulation framework for large, non-stationary spatial data. We observe that for a popular group of statistical models known as spatial autoregressive (SAR) models, the parameters themselves are naturally arranged on a grid [latticekrig, lindgren2011explicit]. Thus, both the spatial fields of interest and their associated parameters can be viewed as images.
Consequently, we adapt image-to-image (I2I) networks—both fully convolutional [ronneberger2015u] and vision transformer based [dosovitskiy2020image, chen2021transunet]—to the task of estimating all model parameters at once. Our networks are chosen to evaluate whether incorporating the attention mechanism [han2022survey, cao2022swin] improves performance over purely convolutional approaches [liu2022convnet], which have far fewer parameters. A key challenge in training these networks is ensuring they recognize complex non-stationarity patterns encountered in large, geoscientific data. Since we seek estimates for images (fields) of statistical parameters, the large, existing corpus of application-specific data [nguyen2023climax, kaltenborn2023climateset, watson2022climatebench] is not appropriate. Instead, we generate our own training data, encoding priors that represent the kind of non-stationary spatial processes expected for geophysical variables. After training, the networks can rapidly infer parameters from a limited number of observed fields, and use the well-known and computationally efficient LatticeKrig [latticekrig] SAR model to cheaply simulate as many synthetic fields as required.
As an illustrative example, we employ the LatticeVision framework on statistics derived from an Earth System Model (ESM). ESM simulations (‘runs’) model the long-term evolution of our climate, providing scientists, policymakers, agricultural planners, and other key stakeholders with critical projections that inform decision making and resilience planning. Additionally, many downstream analyses, such as those that estimate crop yield or flood risk, rely directly on the results of these runs as inputs [rosenzweig2014assessing, wagener2022evaluation, ghosh2009climate]. The computational cost of performing more than a handful of ESM runs is extremely prohibitive, limiting ensemble sizes to 3-100 members [kay2015community, rodgers2021ubiquity]; too few for robust uncertainty quantification of many climate impact models [schwarzwald2022importance], quantifying internal ESM variability [milinski2020large, deser2020insights], or capturing rare events [eyring2024pushing]. Following parameter estimation, we generate realistic ensembles containing thousands of fields in a matter of seconds; a stark contrast to the tens of millions of core-hours required by ESMs.
@@
In experiments with both simulated data and ESM outputs, I2I networks outperform local methods in accuracy—especially for detecting long-range, anisotropic correlations—as well as in computational efficiency. This advantage stems from their ability to estimate parameters for the entire spatial domain simultaneously, rather than by section [sainsbury2024neural] or by pixel [wiens2020modeling]. Our approach also demonstrates improved robustness when inferring weakly-identifiable parameters from a small number of replicate spatial fields, which are represented as channels of the input image.
In summary, our work makes the following contributions:
@@
  • We propose a global framework for efficiently estimating non-stationary parameters with I2I networks, addressing the limitations of existing local approaches.
@@
  • We pair our novel, global estimators with a computationally efficient and flexible SAR model [latticekrig, LKrigpackage] for rapid emulation, and validate this framework on ESM outputs.
All accompanying code for the experiments and results in this paper can be found at github.com/antonyxsik/LatticeVision.
@@ 2 Background: Gaussian Processes and SAR models
The past thirty years have seen the development of statistical models for spatial data that are invaluable for spatial prediction and emulation [heaton2019case, katzfuss2017multi, fuentes2002spectral, latticekrig, frk, guhaniyogi2018meta]. Despite recent advances in generative networks [rombach2022high, ruhling2024probablistic, price2023gencast, li2023seeds], statistical models remain useful when data is limited, and also contain interpretable parameters. These models are typically built around Gaussian processes (GPs). With locations $\mathbf{s}\in\mathbb{R}^{2}$, a mean-zero GP $f(\mathbf{s})$ is fully specified by its covariance function $k(\mathbf{s},\mathbf{s}^{\prime})=\mathbb{E}\left[f(\mathbf{s})f(\mathbf{s}^{\prime})\right]$. The kernel $k$ must be positive definite and this requirement is typically enforced by assuming stationarity and isotropy [cressie2015statistics]. That is, $k$ is independent of the location in the domain (stationarity), and depends only on the separation distance (isotropy):
@@
Here $\mathcal{C}(0)=1$, $\sigma^{2}$ is the variance of the GP, and $\kappa_{m}>0$ governs the spatial correlation range. A popular choice for $\mathcal{C}$ is the Matérn family [^1] In reviewing commonly used covariance kernels, [stein1999interpolation] concludes: ‘Use the Matérn’ [lindgren2011explicit] with an additional parameter, $\nu>0$, that controls the smoothness of process realizations. If $f(\mathbf{s})$ is a realization from the Matérn family with parameter $\nu$, it will have $\lceil\nu\rceil-1$ derivatives in a mean square sense. The Matérn class includes several common kernels as special cases (e.g., exponential and Gaussian), and its parameters have useful physical interpretations in geoscience. The primary obstacle to working with GP-based models is the computational cost ($\mathcal{O}(n^{3})$ operations and $\mathcal{O}(n^{2})$ memory for $n$ spatial locations) of factorizing the covariance matrix $\Sigma\in\mathbb{R}^{n\times n}$ where $\Sigma_{ij}=k(\mathbf{s}_{i},\mathbf{s}_{j})$ [sungeostat].
The SPDE method.
GPs from the Matérn family have an equivalent representation in terms of a stochastic partial differential equation (SPDE) [matern2013spatial, whittle1954stationary]:
@@
where $\kappa^{2}$ controls the correlation range, and is similar but not identical to $\kappa_{m}$ in the Matérn, $\Delta$ is the Laplacian operator, and $\mathcal{W}(\mathbf{s})$ is a white noise Gaussian process with zero mean and variance $\sigma^{2}$. The so-called (SPDE) method [lindgren2011explicit, lindgren2022spde] connects Matérn GPs to Gaussian Markov Random Fields (GMRFs) by demonstrating that discretizing Equation 2 on a regular grid yields a GMRF which approximates the Matérn GP well. This approach can be extended [^2] For simplicity, we shall henceforth focus on the $\nu=1$ case, the so-called Whittle covariance [whittle1954stationary] to incorporate non-stationarity and anisotropy---a necessity for modeling complex geophysical processes such as the Earth’s climate---by incorporating a spatially varying dispersion matrix $H(\mathbf{s})\in\mathbb{R}^{2\times 2}$ and allowing $\kappa^{2}$ to vary in space:
$(\kappa^{2}(\mathbf{s})-\nabla\cdot H(\mathbf{s})\nabla)f(\mathbf{s})=\mathcal{W}(\mathbf{s}).$ (3)
We construct $H(\mathbf{s})$ via its eigendecomposition [^3] This particular form of the eigendecomposition is for consistency with the literature [haskard2007anisotropic]: $H(\mathbf{s})=R(\mathbf{s})^{\top}D(\mathbf{s})^{2}R(\mathbf{s})$ where
$R(\mathbf{s})=\left[\begin{array}[]{cc}\cos\left(\theta(\mathbf{s})\right)&-\sin\left(\theta(\mathbf{s})\right)\\ \sin\left(\theta(\mathbf{s})\right)&\cos\left(\theta(\mathbf{s})\right)\end{array}\right],\quad D(\mathbf{s})=\left[\begin{array}[]{cc}\sqrt{\rho(\mathbf{s})}&0\\ 0&1/\sqrt{\rho(\mathbf{s})}\end{array}\right]$ (4)
The generalized Laplacian in Equation 3 can be written explicitly as
$\nabla\cdot H(\mathbf{s})\nabla\equiv H_{1,1}(\mathbf{s})\frac{\partial^{2}}{\partial x^{2}}+2H_{2,1}(\mathbf{s})\frac{\partial^{2}}{\partial x\partial y}+H_{2,2}(\mathbf{s})\frac{\partial^{2}}{\partial y^{2}}.$ (5)
Interpreting the parameter fields.
@@
By specifying the ‘‘parameter fields’’ $\kappa^{2}(\mathbf{s}),\theta(\mathbf{s})$, and $\rho(\mathbf{s})$ we obtain a rich class of nonstationary GPs. Because we are defining this model in terms of (3), the problem of explicitly specifying an analytical form for the covariance that is positive definite is avoided. Moreover, these parameter fields are interpretable and can yield physical insights into the spatial dependence of the field. Specifically, $\kappa^{2}$ controls the overall range of correlation (larger $\kappa^{2}$ means more localized dependence), $\rho$ controls the degree of anisotropy ($\rho=1$ corresponds to isotropy), and $\theta$ controls the direction of anisotropy; see Figure 2.
Discretizing the SPDE.
@@
To obtain a computable model from (3), one approximates this SPDE using either a finite element or finite difference method. Following [wiens2020modeling] [^4] We correct a minor error in their derivation. Our derivation can be found in Appendix A.3 we use the finite difference method on a regular, evenly spaced grid, yielding the stencil:
$\begin{array}[]{c|c|c}\frac{H_{1,2}(\mathbf{s})}{2}&-H_{2,2}(\mathbf{s})&\frac{-H_{1,2}(\mathbf{s})}{2}\\ \hline\cr-H_{1,1}(\mathbf{s})&\kappa^{2}(\mathbf{s})+2H_{1,1}(\mathbf{s})+2H_{2,2}(\mathbf{s})&-H_{1,1}(\mathbf{s})\\ \hline\cr\frac{-H_{1,2}(\mathbf{s})}{2}&-H_{2,2}(\mathbf{s})&\frac{H_{1,2}(\mathbf{s})}{2}.\end{array}$ (6)
@@
Let $\mathbf{y}\in\mathbb{R}^{n}$ denote a discretized solution to (3) on this grid, with $y_{i,j}$ denoting the value at grid location $(i,j)$. Then $\mathbf{y}$ is the solution to $B\mathbf{y}=\mathbf{e}$ where $\mathbf{e}\sim\mathcal{N}(\mathbf{0},I)$ is a sample from the standard multivariate normal distribution and $B\in\mathbb{R}^{n\times n}$ is the spatial autoregressive (SAR) matrix associated to the stencil (6). Lindgren et al [lindgren2011explicit] show that $\mathbf{y}$ approximates a sample from the Matérn GP associated to (3). As $B$ is sparse and structured (i.e., it is a banded matrix) this linear system may be solved at a computational cost $\mathcal{O}(n^{3/2})$, thus sidestepping the bottleneck associated with working with the GP directly. This formulation results in a GMRF with a particular SAR structure, where, by linear statistics, the precision matrix is computed as $Q=B^{\top}B$.
@@ 3 Non-stationary data generation
In order for the I2I networks to be successful, the training data needs to encode priors appropriate for geoscientific applications. Previous work [wiens2020modeling] has shown that coastlines, long-range East-West correlations due to jet streams, and oceanic circulation yield parameter fields that are important yet challenging to detect. So, we construct a pipeline for generating synthetic fields exhibiting these key phenomena. First, we construct spatially varying parameter fields $\kappa^{2}(\mathbf{s}),\rho(\mathbf{s}),\theta(\mathbf{s})$ and then use them to produce an $M$-replicate ensemble of synthetic fields $Y=\{\mathbf{y}^{(m)}\}^{M}_{m=1}$, where each $\mathbf{y}^{(m)}\in\mathbb{R}^{H\times W}$. The parameter fields are concatenated along the channel dimension to form a three channel, ground truth ‘‘image’’ $\Phi\in\mathbb{R}^{3\times H\times W}$, where $\Phi(\mathbf{s})=[\kappa^{2}(\mathbf{s}),\rho(\mathbf{s}),\theta(\mathbf{s})]^{T}$. The replicates of the synthetic fields are also concatenated along the channel dimension to form the input image to our networks $Y\in\mathbb{R}^{M\times H\times W}$. Thus, we have input-output pairs $(Y,\Phi)$ in our data.
@@
Each time we construct a single parameter field, we first sample one of eight spatial patterns $p^{(i)}(\mathbf{s};\boldsymbol{\Omega}^{(i)})$, $i=1,2,\dots,8$. The patterns are simple spatial functions that dictate how a parameter will vary across the domain, and are designed to be caricatures of real geophysical variability. We hypothesize that these patterns will serve as ‘‘building blocks’’, enabling our networks to generalize to the kinds of parameter changes that are present in geophysical settings. Each pattern is defined by its hyperparameters $\boldsymbol{\Omega}^{(i)}$, which are randomly sampled from a series of prior uniform distributions. For example, when a ‘‘Coastline’’ pattern is selected, values that dictate the position and variation of the ‘‘Coastline’’ are chosen as well. The relative frequency and qualitative descriptions of these patterns are illustrated in Figure 3, with detailed functional forms and hyperparameter priors provided in Appendix B.
Figure 3: Spatial patterns (left) and their frequency (right).
@@
Once a pattern $p^{(i)}$ and its hyperparameters $\boldsymbol{\Omega}^{(i)}$ have been chosen, we then sample values from the prior distribution of the specific parameter $(\kappa^{2},\rho,\text{or }\theta)$ that we are constructing a field for. Each pattern (except ‘‘Constant’’) requires sampling two values which dictate the maximum and minimum value across the resulting parameter field. We set uniform priors on the anisotropy parameters $\rho\sim\mathcal{U}(1,7)$ and $\theta\sim\mathcal{U}\left(-\frac{\pi}{2},\frac{\pi}{2}\right)$, and set a mixed prior of $\kappa^{2}\sim 0.6\log\mathcal{U}(10^{-4},2)+0.4\mathcal{U}(10^{-4},2)$ for the correlation range. These choices ensure we capture a broad range of spatial relationships, ranging from isotropy ($\rho=1$) to very elongated ellipses ($\rho\gg 1$) in any direction.
@@
CNN data.
For the local estimation setting, we assume local stationarity: within a small spatial window, all grid cells are governed by the same parameters. This makes data generation for the CNN comparatively straightforward. Rather than constructing parameter fields, we independently sample a single value for each of $(\kappa^{2},\rho,\theta)$ from the prior distributions defined above. These parameters are then encoded into a SAR matrix $B$, resulting in stationary synthetic fields. We generate smaller fields than for the I2I networks, with each sample consisting of $M$ replicates $\mathcal{Y}\in\mathbb{R}^{M\times h\times w}$, where $h\ll H$ and $w\ll W$. The associated ground truth is a parameter vector $\phi=[\kappa^{2},\rho,\theta]^{T}\in\mathbb{R}^{3}$. Thus, our input-output pairs are $(\mathcal{Y},\phi)$.

### v2

@@ Abstract
In many applications, we wish to fit a parametric statistical model to a small ensemble of spatially distributed random variables (‘fields’). However, parameter inference using maximum likelihood estimation (MLE) is computationally prohibitive, especially for large, non-stationary fields. Thus, many recent works train neural networks to estimate parameters given spatial fields as input, sidestepping MLE completely. In this work we focus on a popular class of parametric, spatially autoregressive (SAR) models. We make a simple yet impactful observation; because the SAR parameters can be arranged on a regular grid, both inputs (spatial fields) and outputs (model parameters) can be viewed as images. Using this insight, we demonstrate that image-to-image (I2I) networks enable faster and more accurate parameter estimation for a class of non-stationary SAR models with unprecedented complexity.
@@ 1 INTRODUCTION
Modeling large, gridded spatial data has become a central challenge in many scientific and industrial applications. This typically involves fitting parametric spatial models to enable prediction, data fusion, and, when data are limited, rapid simulation of additional fields. The bottleneck in this framework is inferring the parameters of the statistical model using maximum likelihood estimation (MLE), which becomes computationally intractable as dataset size increases [73, 75]. Moreover, spatial data over large domains frequently exhibit non-stationarity, meaning key parameters vary over space, further complicating parameter estimation. Many recent works have replaced MLE with neural networks, mapping spatial fields directly to local parameter estimates [38, 1, 87, 34]. Like MLE, these methods are limited to dividing large fields into smaller sections, each assumed to be stationary, and estimating parameters independently for each section. Although faster than MLE, the number of forward passes required scales linearly with the number of sections. Additionally, such local neural estimators struggle to capture long-range, global context.
@@
Figure 1: An illustration of the LatticeVision workflow applied to ESM outputs. Spatial fields are fed into an I2I network, which in turn produces estimates of the non-stationary parameter fields. These are encoded into a SAR model from which synthetic replicates are efficiently simulated.
@@
In this work we introduce LatticeVision, a global estimation and emulation framework for large, non-stationary spatial data. We observe that for a popular group of statistical models known as spatial autoregressive (SAR) models, the parameters themselves are naturally arranged on a grid. Thus, both the spatial fields of interest and their associated parameters can be viewed as images. Consequently, we adapt image-to-image (I2I) networks—both fully convolutional [60] and vision transformer based [16, 10]—to the parameter estimation task. Unlike local neural estimators, I2I networks estimate all model parameters at once, in a single forward pass. Our networks are chosen to evaluate whether incorporating the attention mechanism improves performance over purely convolutional approaches [40], which have far fewer parameters. We find that a hybrid approach offers the best performance.
A key challenge in training these networks is ensuring they recognize complex non-stationarity patterns encountered in large, geoscientific data. Since we require (field, parameters) pairs for training, the existing corpora of application specific fields [46, 25, 79] are inapplicable. Thus we generate our own training data, encoding priors that represent the kind of non-stationary spatial processes expected for geophysical variables.
@@
In experiments with simulated data, I2I networks outperform local neural estimators in both speed and accuracy. This advantage stems from the fact that I2I networks process the entire field at once, rather than by section [65] or by pixel [82]. We also show that I2I networks estimate weakly-identifiable parameters from a small number of replicate spatial fields more reliably than previous, local approaches.
As an illustrative example, we employ the LatticeVision framework on data from multiple Earth System Models (ESMs). ESM simulations (‘runs’) model the long-term evolution of Earth’s climate, providing decision-makers with critical projections. The computational cost of performing more than a handful of ESM runs is extremely prohibitive, limiting ensemble sizes to 3-100 members [28, 59]; too few for many applications [44, 15, 66, 18]. Following parameter estimation, we generate realistic ensembles containing thousands of fields in a matter of seconds using the LatticeKrig package [48]; a stark contrast to the tens of millions of core-hours required by ESMs. We show that ensembles simulated with parameters estimated by I2I networks better capture spatial relationships—especially long-range anisotropic correlations—than those produced by local methods.
In summary, we make the following contributions:
@@
  • We propose a global framework for efficiently estimating non-stationary parameters with I2I networks, demonstrate that hybrid architectures outperform those that are purely convolutional or transformer-based for this task, and address the limitations of existing local approaches.
@@
  • We pair our novel, global estimators with a computationally efficient and flexible SAR model [48, 50], and validate this framework on ESM outputs.
All code is submitted as supplementary material, and will be released following the review process.
@@ 2 BACKGROUND: GAUSSIAN PROCESSES AND SAR MODELS
The past thirty years have seen the development of statistical models for spatial data that are invaluable for spatial prediction and emulation [23, 27, 19, 21]. Despite recent advances in generative deep learning for spatial data [62, 55, 35], statistical models outperform other approaches when data is limited [42]. Moreover, statistical models contain interpretable parameters, which are useful for downstream applications.
We consider statistical models built around Gaussian processes (GPs). With locations $\mathbf{s}\in\mathbb{R}^{2}$, a mean-zero GP $f(\mathbf{s})$ is fully specified by its covariance function $k(\mathbf{s},\mathbf{s}^{\prime})=\mathbb{E}\left[f(\mathbf{s})f(\mathbf{s}^{\prime})\right]$. The kernel $k$ must be positive definite and this requirement is typically enforced by assuming stationarity and isotropy [14]. That is, $k$ is independent of the location in the domain (stationarity), and depends only on the separation distance (isotropy):
@@
Here $\mathcal{C}(0)=1$, $\sigma^{2}$ is the variance of the GP, and $\kappa_{m}>0$ governs the spatial correlation range. A popular choice for $\mathcal{C}$ is the Matérn family [^1] In reviewing commonly used covariance kernels, [72] concludes: ‘Use the Matérn’. with an additional smoothness parameter, $\nu>0$. The Matérn class includes several common kernels as special cases (e.g., exponential and Gaussian). The primary obstacle to working with GP-based models is the computational cost ($\mathcal{O}(n^{3})$ operations and $\mathcal{O}(n^{2})$ memory for $n$ spatial locations) of factorizing the covariance matrix $\Sigma\in\mathbb{R}^{n\times n}$ where $\Sigma_{ij}=k(\mathbf{s}_{i},\mathbf{s}_{j})$ [75].
The SPDE Method
GPs from the Matérn family have an equivalent representation in terms of a stochastic partial differential equation (SPDE) [43, 81]:
@@
where $\kappa^{2}$ controls the correlation range, and is similar but not identical to $\kappa_{m}$ in the Matérn, $\Delta$ is the Laplacian operator, and $\mathcal{W}(\mathbf{s})$ is a white noise Gaussian process with zero mean and variance $\sigma^{2}$. The so-called SPDE method [37, 36] connects Matérn GPs to Gaussian Markov Random Fields (GMRFs) by demonstrating that discretizing Equation (2) on a regular grid yields a GMRF which approximates the Matérn GP well. This approach can be extended [^2] For simplicity, we shall henceforth focus on the $\nu=1$ case, the so-called Whittle covariance [81] to allow non-stationarity and anisotropy by incorporating a spatially varying dispersion matrix $D(\mathbf{s})\in\mathbb{R}^{2\times 2}$ and letting $\kappa^{2}$ vary in space:
$(\kappa^{2}(\mathbf{s})-\nabla\cdot D(\mathbf{s})\nabla)f(\mathbf{s})=\mathcal{W}(\mathbf{s}).$ (3)
Following [22], we construct $D(\mathbf{s})$ via its eigendecomposition: $D(\mathbf{s})=R(\mathbf{s})^{\top}\Lambda(\mathbf{s})R(\mathbf{s})$ where
$R(\mathbf{s})=\begin{bmatrix}\cos\theta(\mathbf{s})&-\sin\theta(\mathbf{s})\\ \sin\theta(\mathbf{s})&\cos\theta(\mathbf{s})\end{bmatrix},\\ \Lambda(\mathbf{s})=\begin{bmatrix}\rho(\mathbf{s})&0\\ 0&\frac{1}{\rho(\mathbf{s})}\end{bmatrix}$ (4)
The generalized Laplacian in (3) can be written as
$\nabla\cdot D(\mathbf{s})\nabla\equiv D_{1,1}(\mathbf{s})\frac{\partial^{2}}{\partial x^{2}}+2D_{2,1}(\mathbf{s})\frac{\partial^{2}}{\partial x\partial y}+D_{2,2}(\mathbf{s})\frac{\partial^{2}}{\partial y^{2}}.$ (5)
Interpreting the Parameter Fields
@@
By specifying the “parameter fields” $\kappa^{2}(\mathbf{s}),\theta(\mathbf{s})$, and $\rho(\mathbf{s})$ we obtain a rich class of nonstationary GPs. Because we are defining this model in terms of (3), the problem of explicitly specifying an analytical form for the covariance that is positive definite is avoided. Moreover, these parameter fields are interpretable and can yield physical insights into the spatial dependence of the field. Specifically, $\kappa^{2}$ controls the overall range of correlation (larger $\kappa^{2}$ means more localized dependence), $\rho$ controls the degree of anisotropy ($\rho=1$ corresponds to isotropy), and $\theta$ controls the direction of anisotropy; see Figure 2.
Discretizing the SPDE
@@
To obtain a computable model from (3), one approximates this SPDE using either a finite element or finite difference method. Following [82] [^3] We correct a minor error in their derivation. Our derivation can be found in Appendix A.3 we use the finite difference method on a regular grid, yielding the stencil:
$\begin{array}[]{c|c|c}\frac{D_{1,2}(\mathbf{s})}{2}&-D_{2,2}(\mathbf{s})&\frac{-D_{1,2}(\mathbf{s})}{2}\\ \hline\cr-D_{1,1}(\mathbf{s})&\kappa^{2}(\mathbf{s})+2D_{1,1}(\mathbf{s})+2D_{2,2}(\mathbf{s})&-D_{1,1}(\mathbf{s})\\ \hline\cr\frac{-D_{1,2}(\mathbf{s})}{2}&-D_{2,2}(\mathbf{s})&\frac{D_{1,2}(\mathbf{s})}{2}.\end{array}$ (6)
@@
Let $\mathbf{y}\in\mathbb{R}^{n}$ denote a discretized solution to (3) on this grid, with $y_{i,j}$ denoting the value at grid location $(i,j)$. Then $\mathbf{y}$ is the solution to $B\mathbf{y}=\mathbf{e}$ where $\mathbf{e}\sim\mathcal{N}(\mathbf{0},I)$ is a sample from the standard multivariate normal distribution and $B\in\mathbb{R}^{n\times n}$ is the spatial autoregressive (SAR) matrix associated to the stencil (6). [37] show that $\mathbf{y}$ approximates a sample from the Matérn GP associated to (3). As $B$ is sparse and structured (i.e., it is a banded matrix) this linear system may be solved at a computational cost $\mathcal{O}(n^{3/2})$, thus sidestepping the bottleneck associated with working with the GP directly. This formulation results in a GMRF with a particular SAR structure, where, by linear statistics, the precision matrix is $Q=B^{\top}B$.
@@ 3 NON-STATIONARY DATA GENERATION
In order for the I2I networks to be successful, the training data needs to encode priors appropriate for geoscientific applications. Previous work [82] has shown that coastlines, long-range East-West correlations due to jet streams, and oceanic circulation yield parameter fields that are important yet challenging to detect. So, we construct a pipeline for generating synthetic fields exhibiting these key phenomena. First, we construct spatially varying parameter fields $\kappa^{2}(\mathbf{s}),\rho(\mathbf{s}),\theta(\mathbf{s})$ and then use them to produce an $M$-replicate ensemble of synthetic fields $Y=\{\mathbf{y}^{(m)}\}^{M}_{m=1}$, where each $\mathbf{y}^{(m)}\in\mathbb{R}^{H\times W}$ is generated using the same parameter fields. These parameter fields are concatenated along the channel dimension to form a three channel, ground truth “image” $\Phi\in\mathbb{R}^{3\times H\times W}$, where $\Phi(\mathbf{s})=[\kappa^{2}(\mathbf{s}),\rho(\mathbf{s}),\theta(\mathbf{s})]^{T}$. The replicates of the synthetic fields are also concatenated along the channel dimension to form the input image to our networks $Y\in\mathbb{R}^{M\times H\times W}$. Thus, we have input-output pairs $(Y,\Phi)$ in our data.
@@
Each time we construct a single parameter field, we first sample one of eight spatial patterns $p^{(i)}(\mathbf{s};\boldsymbol{\Omega}^{(i)})$, $i=1,2,\dots,8$. The patterns are simple spatial functions that dictate how a parameter will vary across the domain, and are designed to be caricatures of real geophysical variability. We hypothesize that these patterns will serve as “building blocks”, enabling our networks to generalize to the kinds of parameter changes present in geophysical settings. Each pattern is defined by its hyperparameters $\boldsymbol{\Omega}^{(i)}$, which are randomly sampled from a series of prior uniform distributions. For example, when a “Coastline” pattern is selected, values that dictate the position and variation of the “Coastline” are chosen as well. The relative frequency and qualitative descriptions of these patterns are illustrated in Figure 3, with detailed functional forms and hyperparameter priors provided in Appendix B.
Figure 3: Spatial patterns and their frequencies.
@@
Once a pattern $p^{(i)}$ and its hyperparameters $\boldsymbol{\Omega}^{(i)}$ have been chosen, we sample values from the prior distribution of the specific parameter $(\kappa^{2},\rho,\text{or }\theta)$ for which we are constructing a field. Each pattern (except “Constant”) requires sampling two values which dictate the maximum and minimum value across the resulting parameter field. We set uniform priors on the anisotropy parameters $\rho\sim\mathcal{U}(1,7)$ and $\theta\sim\mathcal{U}\left(-\frac{\pi}{2},\frac{\pi}{2}\right)$, and set a mixed prior of $\kappa^{2}\sim 0.6\log\mathcal{U}(10^{-4},2)+0.4\mathcal{U}(10^{-4},2)$ for the correlation range. These choices ensure we capture a broad range of spatial relationships, ranging from isotropy ($\rho=1$) to very elongated ellipses ($\rho\gg 1$) in any direction.
@@
CNN Data
For the local estimation setting, we assume local stationarity: within a small spatial window, all grid cells are governed by the same parameters. This makes data generation for the local CNNs comparatively straightforward. Rather than constructing parameter fields, we independently sample a single value for each of $(\kappa^{2},\rho,\theta)$ from the prior distributions defined above. These parameters are then encoded into a SAR matrix $B$, resulting in stationary synthetic fields. We generate smaller fields than for the I2I networks, with each sample consisting of $M$ replicates $\mathcal{Y}\in\mathbb{R}^{M\times h\times w}$, where $h\ll H$ and $w\ll W$. The associated ground truth is a parameter vector $\phi=[\kappa^{2},\rho,\theta]^{T}\in\mathbb{R}^{3}$. Thus, our input-output pairs are $(\mathcal{Y},\phi)$.